Find the sum of the vector using the component method. Please show the solution.



(vec) A = 5m, 25°


(vec) B = 11m, 125°


(vec) C = 25m, 35°



Note:


(vec) R = (vec) A + (vec) B + (vec) C

1
Expert's answer
2021-11-16T10:01:13-0500

Find the components:


"A_x=5\cos25°=4.5\text{ m},\\
A_y=5\sin25°=2.11\text{ m}.\\
B_x=11\cos125°=-6.31\text{ m},\\
B_y=11\sin125°=9.01\text{ m}.\\
C_x=25\cos35°=20.4\text{ m},\\
C_y=25\sin35°=14.3\text{ m}.\\\space\\
R_x=A_x+B_x+C_x=18.6\text{ m},\\
R_y=A_y+B_y+C_y=25.4\text{ m}.\\\space\\
R=\sqrt{R_x^2+R_y^2}=27.9\text{ m}.\\\space\\
\theta=\arctan\frac{R_y}{R_x}=53.8°."